Intergate completely.

∫(4x^5-15x^4-x^3+12x^2+2x+3)/(x^2-3x-2)dx


1
Expert's answer
2021-02-24T07:04:25-0500

Answer is

"x^{4}-x^{3}-x^{2}+\ln \left|\left(x-\frac{3}{2}\right)^{2}+\frac{1}{4}\right|+C"

Explanation below